OpenStudy (anonymous):

what is the derivative of x^2 sin 8x

OpenStudy (anonymous):

2xsin8x+8x^2cos8x

OpenStudy (campbell_st):

use the product rule...

OpenStudy (anonymous):

Where does the 8x^2 in the second part of the problem come from?

OpenStudy (anonymous):

dsin8x is 8 cos8x multiplied by x^2 is second term

OpenStudy (anonymous):

Okay, I understand now. Thank you for your time. :)

OpenStudy (anonymous):

dsin8x= cos8x d8x=cos8x 8 dx

OpenStudy (anonymous):

\[h(x) = f(x)g(x)\\h'(x) = f'(x)g(x)+f(x)g'(x)\]

OpenStudy (campbell_st):

start with the 2 functions that are the factors of the whole function \[f(x) = x^2...... then....... f'(x) = 2x\] and \[g(x) = \sin(8x).... then..... g'(x) = 8\cos(8x)\] so the product rule says y' =f(x)*g'(x) + g(x)*f'(x) hope this helps